The height of a right triangle is 3 feet more than three times the length of its base. Express the area of the triangle as a function of the length of its base, x, in feet.
Answer by mananth(16946)

base =x
height = 3x+3
A= 1/2 base * h
A= 1/2 * x*(3x+3)
= 3/2*x(x+1)